OpenStudy (anonymous):

how to solve x^2-x-6/ x^2-4 by hand?

OpenStudy (anonymous):

what do you mean "solve"? maybe "factor and cancel"?

OpenStudy (anonymous):

Express x^2-4 as difference of ---->two<--- squares

OpenStudy (anonymous):

\[\frac{x^2-x-6}{x^2-4}=\frac{(x-3)(x+2)}{(x-2)(x+2)}\] is a start